The demand and supply for a one year discount bond with a face value of $1000 is given by Bd: P = -0.5Qd + 1250 and Bs: P = Qs - 250. What is the equilibrium price and quantity for this discount bond? What is the interest rate?
Answer #1

At equilibrium,

-.5*Q + 1250 = Q - 250

1.5*Q = 1500

Q = 1500/1.5

Q = 1000

Price = -.5*1000 + 1250

Price = $750

So, equilibrium price is  $750 and quantity is 1000.

Let, interest rate = R

750 = 1000/(1+R)

R = 1000/750 - 1

R = 33.33%

So, the interest rate is 33.33%